Using the recent monthly expenses and revenue as your baseline, develop a revenue and expense forecast for July - September. Assume there is a direct relationship between Revenue/Expenses and Effective Business Days.

Revenue Expenses Effective Business Days
Apr-08        4,000         3,200 20.0
May-08        4,500         3,650 20.5
Jun-08        6,200         4,300 22.0
Jul-08        6,200 19.0
Aug-08 22.5
Sep-08 21.5
Average Daily $       235 $        178
Answer #1

Effectivee Revenue Expenses Business days Apr-08 May-08 Jun-08 Jul-08 Aug-08 Sep-08 4,000.003,200.00 4,500.00 3,650.00 6,200.

I assume that, there is a direct relationship between Revenue/Expenses and Effective Business Days.

Total of effective working days = 125.50

Daily average revenue = 235

Daily average expenses = 178

So, we take the grand total of total revenue by multiplying

daily average revenue with total effective business days = 235*125.50 = 29492.50

Total revenue of first four months = 4000+4500+6200+6200 = 20,900/-

Revenue of last two months = 29492.50 - 20900 = 8592.50

Last two months revenue allocated on proportionate basis (based on number of effective business days)

Total of effective working days = 125.50

total of effective working days of first four months = 20+20.5+22+19 = 81.5

Balance effective hours(two months) = 125.5 - 81.5 = 44

Therefore,

Revenue for august = 8592.50 *22.50/44

= 4393.89

Revenue for september = 8592.50 * 21.50/44

=4198.61

Expenses for the month of july, august and september is also calculated on proportionate basis

Total of effective working days = 125.50

Daily average expenses = 178

So, we take the grand total of total revenue by multiplying

daily average expenses with total effective business days = 178*125.50 = 22339

Total expenses for first 3 months(april, May, june) = 3200+3650+4300 = 11150

and

Expenses of last three months = 22339 - 11150 = 11189

Last three months expenses allocated on proportionate basis (based on number of effective business days)

Total of effective working days = 125.50

Total of effective working days of first three months = 20+20.5+22 = 62.50

Balance effective hours(three months) = 125.50 - 62.50 = 63

Therefore,

Expenses for  july = 11189 * 19/63

=3374.46

Expenses for august = 11189 * 22.5/63

= 3996.07

Expenses for September = 11189 * 21.50/63

= 3818.47